Hello, this problem involves with plugging values into a shape formula of a trapezoid. I believe the book had given me a wrong answer. Could someone please help validate my answer. I would really appreciate, thank you.

Here is the trapezoid formula:
A = 1/2h(b+B)
Values given are:
A = 91, h = 7 and b = 12

Answer by rothauserc(4718)

A = 1/2h(b+B)
Values given are:
A = 91, h = 7 and b = 12
*******************************
I assume you need to solve for B
********************************
A = 1/2h(b+B)
multiply both sides of = by 2
2A = h(b+B)
divide both sides of = by h
b+B = (2A/h)
subtract b from both sides of =
B = (2A/h) - b
B = ((2*91) / 7) -12
B = 26 - 12 = 14
B = 14